OpenStudy (anonymous):

\[\int\limits {1 \over x \ln x}dx----->x=e^u\]

OpenStudy (anonymous):

Step by step pleasee

OpenStudy (experimentx):

let ln x = u

OpenStudy (anonymous):

Put u = ln x and du = 1/x dx

OpenStudy (experimentx):

du = dx/x

OpenStudy (anonymous):

You have to use x= e^u

OpenStudy (experimentx):

that is in the later part ... for now do above

OpenStudy (anonymous):

But it says you have to substitute first, with x.

OpenStudy (experimentx):

find the value of integration in u, and put the value of u=e^x and you will have your answer.

OpenStudy (anonymous):

I don't know what e^u ln e^u becomes

OpenStudy (anonymous):

ln e^u = u assuming uis +

OpenStudy (anonymous):

\[Isn't~\it~~ \int\limits {1 \over xlnx}dx ---- > \int\limits {1 \over e^u \ln e^u}dx?\]

OpenStudy (experimentx):

no 

OpenStudy (anonymous):

But I have to substitute the x's as e^u

OpenStudy (anonymous):

You should get: \[\int\limits 1/\ln e^{u} du\]

OpenStudy (anonymous):

@order x = e^u implies that dx= e^u du.So the e^u in the denominator cancels out.

OpenStudy (anonymous):

I don't think it means dx =e^u. I think it implies x=e^u

OpenStudy (anonymous):

If x= e^u then dx/du = e^u implies dx =e^u du

OpenStudy (anonymous):

So, can you start from the beginning beginning?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

\[\int\limits {1 \over xlnx}dx---> becomes??\]

OpenStudy (anonymous):

\[\int\limits 1/\ln e^{u} du\]

OpenStudy (anonymous):

\[So,~ what~happens~with~the~first~x?\]